Beta​ Laundry's cost function is

​C(q)equals=70+24q+q^2.

How much does it produce if p=​$60​?

The firm produces

q=______units.

​(Enter your response as a whole​ number.)If the government imposes a per unit tax of t=22​,

what quantity maximizes its​ after-tax profit?

​(Hint: per unit tax increases​ firm's MC and AVC by the amount of​ tax, for instance if per unit tax is 2​ dollars, MC increases by 2 dollars and AVC increases by 2​ dollars)

With the​ tax, the​ profit-maximizing quantity for the firm to produce is

q=_____units. ​(Enter your response as a whole​ number.)

Does it operate or shut​ down?

The firm should produce/shutdown?
Answer #1

a) MC = dC/dq = 24 + 2q

Profit is maximized where P = MC = 60

q = 36/2 = 18 units

b) Per unit tax = 22

MC = 22 + 24 + 2q = 46 + 2q

P = MC :

60 = 46 + 2q

q = 7 units

The firm should produce as it is able to recover its fixed costs. (P > min AVC)